9. Suppose the nominal rate you see on a Treasury bill is 4.31 percent and its real rate is 2.5 percent, what is the inflation rate?

1.77% Please Show All Work
Answer #1

Based on Fischer relation,

(1 + Nominal rate) = (1 + Real rate) * (1 + Inflation)

{In some texts, you would find this relation as Nominal Rate = Real Rate + Inflation. This is an approximate version of the above mentioned actual relation}

(1 + 4.31%) = (1 + 2.5%) * (1 + Inflation)

(1 + Inflation) = 1.0431/1.025 = 1.01766

Inflation = 0.01766

Inflation = 1.77%